OpenStudy (anonymous):

Write a general linear equation for the line through the two points (-2,1), (2,-2). Shall I first find slope and then use the point-slope formula???

OpenStudy (anonymous):

@amistre64 Am I correct or is there some other method to be used @satellite73 Am I correct or is there some other method to be used

OpenStudy (amistre64):

you can if you want to yes

OpenStudy (anonymous):

is there another method??

OpenStudy (amistre64):

or, you can setup 2 equations: y = mx + c y = mx + k and equate them

OpenStudy (amistre64):

well, +c for both of them would be better (-2,1) 1 = -2m + c (2,-2) -2 = 2m + c

OpenStudy (anonymous):

What do I do after this??

OpenStudy (anonymous):

@amistre64 Pls inform what further steps do I have to go through

OpenStudy (amistre64):

well, its either a substitution or elimination to determine m and x

OpenStudy (amistre64):

m and c that is

OpenStudy (anonymous):

Ok, so it is same as solving simultaneous linear equations

OpenStudy (amistre64):

1 = -2m + c -2 = 2m + c ------------ -1 = c 1 = -2m -1 2 = -2m -1 = m the line equation then amounts to: y = -x-1

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thanks for your help....☺

OpenStudy (amistre64):

the general setup tho is to convert that into: Ax + By = C or Ax + By + C = 0

OpenStudy (amistre64):

your original method would have worked well too :) good luck